OpenStudy (anonymous):

Solve for y in the equation 4y+15=-5y+9 a)-6 b) -2/3 c)2/3 d)6

OpenStudy (anonymous):

4y + 15 = -5y + 9 4y + 5y = 9 -15 9y = -6 y = -6/9 y = -2/3

OpenStudy (anonymous):

to solve for "Y" move them all onto one side, and all your non "Y"'s on the other 4y+5y=9-15 now simplify 9y=-6 solve for "Y" y=-6/9 reduce -2/3 is your answer (:

OpenStudy (anonymous):

thank u guys<3 :)